OpenStudy (anonymous):

Sam is observing the velocity of a car at different times. After three hours, the velocity of the car is 51 km/h. After five hours, the velocity of the car is 59 km/h. Part A: Write an equation in two variables in the standard form that can be used to describe the velocity of the car at different times. Show your work and define the variables used. (5 points)

OpenStudy (anonymous):

I TAG u in something it has the answer

OpenStudy (vickyisthesmartone123):

Hold on i know the answer

OpenStudy (vickyisthesmartone123):

ok, to your problem, you have 2 things: hour and velocity, right? it says 3 hours--------51m/s 5hours ---------59m/s got it?

OpenStudy (vickyisthesmartone123):

@Lydia4goals You there?

OpenStudy (anonymous):

i don't think so

OpenStudy (vickyisthesmartone123):

Now, ignore the word, just focus on the numbers, you have 2 points (3, 51) and (5,59) . Got what I mean?

OpenStudy (vickyisthesmartone123):

I'm just goign to explain ti then she can figure it out on her own

OpenStudy (anonymous):

okay yeah

OpenStudy (vickyisthesmartone123):

to write an equation of a line, you need : 1/ slope. (can get it from 2 points) 2/ 1 point .

OpenStudy (vickyisthesmartone123):

ok good your here

OpenStudy (vickyisthesmartone123):

1/ SLOPE: We have a formula to find it from 2 given points, it says m = y2−y1x2−x1 can you show me how to apply our 2 points to this formula? YOUR TURN!!!

OpenStudy (anonymous):

The slope would be 4 right?

OpenStudy (anonymous):

sorry i was trying to solve it

OpenStudy (vickyisthesmartone123):

you are correct, now 2/ pick 1 point, either of them , which one?

OpenStudy (anonymous):

(3, 51).

OpenStudy (vickyisthesmartone123):

ok, the last step is put slope, point into formula: y−y0=m(x−x0) YOUR TURN!!!

OpenStudy (anonymous):

59-51=4(x-3)??

OpenStudy (vickyisthesmartone123):

correct now...

OpenStudy (anonymous):

huh?

OpenStudy (vickyisthesmartone123):

What do you mean huh?

OpenStudy (vickyisthesmartone123):

y = 4(x - 3) + 51 you should get this then open the parentheses.

OpenStudy (vickyisthesmartone123):

so, our equation should be v = 4t +39 THIS IS YOUR ANSWER <--- Your welcome! :)